OpenStudy (anonymous):

if x and y are non real cube roots of unity then x^7 + y^7 = ?

OpenStudy (lgbasallote):

non real polynomial of unity

OpenStudy (anonymous):

non real cube root

OpenStudy (anonymous):

put x as \(\omega^2\) and y as \(\omega^3\)

hartnn (hartnn):

if x^3=1,y^3=1 whats x^7 ? and y^7 ??

hartnn (hartnn):

@Yahoo! it will be x+y , right?

OpenStudy (anonymous):

\[\large (w^2)^7 + (w^3)^7 = = w^{14} + w^{21}\]

OpenStudy (anonymous):

nop....The answer should be -1/xy

hartnn (hartnn):

@waterineyes how is your x w^2 ??

OpenStudy (anonymous):

\[w^3 = 1\]

OpenStudy (anonymous):

Soryyyyyyyyyyyyyyyyyyyyyy

OpenStudy (anonymous):

\[(w)^7 + (w^2)^7 = w^7 + w^{14} = w^7(1 + w^7)\]

OpenStudy (anonymous):

i got answer as -1

OpenStudy (anonymous):

(w^7) = (w^3)^2 *w = w

hartnn (hartnn):

yup, -1 is the answer and -1/xy is same as -1 because xy=1 check it out.

OpenStudy (anonymous):

\[w^3 = 1 \implies w^7 = w^6 \cdot w = w\]

OpenStudy (anonymous):

Yes....)

OpenStudy (anonymous):

\[w^7 + w^{14} = w + w^2\]

OpenStudy (anonymous):

By cube root of unity: \[1 + w + w^2 = 0\] So: \[w + w^2 = -1\]

OpenStudy (anonymous):

thxxxxx